# How Does A Text Preprocessing Pipeline Affect Ontology Syntactic Matching?

### v3

@@ VI-A LLMs Used in OM Text Preprocessing
    Considering the complexity of prompts, we also test an additional case where the prompts have a self-reflection phase (i.e. add “Write a short explanation” to the prompt). We use parentheses to indicate that the different responses generated by the prompts have a self-reflection phase.
Table III shows the results of using LLMs for syntactic matching without text preprocessing. We expect the LLMs to generate “Yes” answers for this TP sample. However, the majority of the LLMs output the “No” answers (marked with the red colour in the table) without text preprocessing.
LLMs <KEYWORD> “identical” “interchangeable” “equivalent” gpt-4-turbo No No No gpt-3.5-turbo No No No claude-3-opus No Yes Yes claude-3-sonnet No No No claude-3-haiku No No No mistral-large No No Yes mistral-medium No No No mistral-small No Yes Yes (No) llama-3-70b No No No llama-3-8b Yes Yes Yes
TABLE III: Using LLMs for syntactic matching without text preprocessing. Prompt: “Is ArtGallery <KEYWORD> to/with art_gallery? Answer yes or no. (Write a short explanation.)”
Table IV shows the results of using LLMs for syntactic matching with text preprocessing. We can see that the error rate can be significantly reduced, but some LLMs still continue to produce incorrect results. The result also shows that LLMs with a larger number of parameters (e.g. llama-3-70b) and prompts with a self-reflection phase (labelled in parentheses) offer “too much of a good thing”. They do not help with syntactic matching and may even produce more errors in OM tasks. Although this preliminary study does not cover all possible triggers used in prompt engineering, it is adequate to highlight the importance of proper preprocessing before fitting the text into the LLMs. We believe that the classical programming-based text preprocessing pipeline is more stable and reliable than using LLM’s implicit text preprocessing or prompt-based text preprocessing for OM tasks.
LLMs <KEYWORD> “identical” “interchangeable” “equivalent” gpt-4-turbo Yes Yes Yes gpt-3.5-turbo Yes Yes Yes claude-3-opus Yes Yes Yes claude-3-sonnet Yes (No) No Yes (No) claude-3-haiku No No Yes (No) mistral-large Yes Yes Yes mistral-medium No Yes Yes mistral-small Yes Yes Yes llama-3-70b Yes Yes (No) Yes (No) llama-3-8b Yes Yes Yes
TABLE IV: Using LLMs for syntactic matching with text preprocessing. Prompt: “Is art gallery <KEYWORD> to/with art gallery? Answer yes or no. (Write a short explanation.)”
VI-B LLMs Used in OM Text Preprocessing Pipeline Repair
Using LLMs for OM text preprocessing is not without merit. With strong background knowledge, they can be used to facilitate several subtasks in OM. For example, LLMs can be used as a repair tool for the text preprocessing pipeline, as an alternative to the approach we proposed in Section V-B. The slight difference is that the context-based pipeline repair approach is an ad hoc repair prior to syntactic matching, while the LLM-based approach is a post hoc repair after syntactic matching. We conduct a preliminary study on LLMs used in OM text preprocessing pipeline repair. The experiment settings are described as follows:
@@ VI-B LLMs Used in OM Text Preprocessing Pipeline Repair
    The experiment is set up on the largebio track repository, a biomedical track repository requires higher Precision with low rates of FPs. The Lancaster Stemmer always shows lower Precision across 6 pairs of alignments, meaning that the results contain a significant number of FPs.
@@
    We choose the same 10 LLMs used in the previous experiment in Appendix VI-A.
@@
    We use the prompt “Is X equivalent to Y?” and preprocess the entity names X and Y with N and T before fitting them into the LLMs. The previous experiment in Appendix VI-A shows these two settings are more suitable for OM tasks.
@@
    Considering the complexity of prompts, we also test an additional case where the prompts have a self-reflection phase (i.e. add “Write a short explanation” to the prompt).
Figure 11: The discovery rate of FPs using LLMs on the largebio track repository.
Fig. 11 shows the discovery rate of FPs using 10 different LLMs. We expect the LLMs to generate “No” answers for these FP samples. The GPT model gpt-4-tubo, the Claude model claude-3-haiku, the Mistral model mistral-medium, and the Llama model llama-3-70b achieve better performance in detecting FPs. For the GPT and Llama series, LLMs with a larger number of parameters perform better than those with a relatively smaller number of parameters. For the Claude and Mistral series, LLMs with larger parameters do not improve performance. In the test cases we analysed, they are even less effective than those with a relatively smaller number of parameters. Unlike detecting TPs, we can see a self-reflection phase improves the overall performance of detecting FPs.
VI-C Implications
LLMs were originally developed for the purpose of question-answering (QA). OM is not only a QA task, but also a knowledge-intensive task that relates to a comprehensive understanding and reasoning of domain knowledge. The studies in this section demonstrate the limitations and opportunities of using LLMs for OM tasks. From our experience, we highlight two pathways to customise LLM for a new domain task. One pathway is function calling (FC), which refers to teaching LLMs using external tools. In LLM-based text preprocessing, we could package the programming-based text preprocessing pipeline as a tool for LLMs to use. The other pathway is fine-tuning (FT), where LLMs are customised with local data. In LLM-based text preprocessing repair, we could use human-in-the-loop to validate the repaired results generated by LLMs and feed the validated data back into LLMs, so that they can better understand the context of the task.
VII Discussion
Ontologies provide formalised conceptualisations of knowledge graphs (KGs). Incorporating ontologies for KG-related NLP tasks can enhance the capability to handle complex concepts and relations, thereby improving the quality of KGs built from text. Ontologies also assist interoperability of text generated from KGs. However, aligning and integrating heterogeneous ontologies remains a challenge when using them for NLP tasks. OM aims to bridge the gap by capturing similar concepts that occur in different ontologies.
OM is a complex process that combines syntactic, lexical, and semantic matching. Each level of matching may use different matching techniques, for example, text preprocessing for syntactic matching, text vectorisation for lexical matching, and logical reasoning for semantic matching. This paper concentrates on text preprocessing, a common practice in syntactic matching. While generic text preprocessing pipeline methods in syntactic OM are usually applied on the basis of intuition or extrapolation from other experience, this work advances the state of knowledge towards making design decisions objective and supported by evidence.
@@ VII Discussion
    Our study shows (i) whether to use, or not use, and (ii) how to use these text preprocessing methods. Such experimental results will benefit decision making when selecting appropriate text preprocessing methods for syntactic OM. For large-scale OM, it can significantly reduce unnecessary trial costs.
  2. (2)
    Our context-based pipeline repair approach is proposed to repair the less effective Phase 2 text preprocessing methods. Its broader value is showing how to maximise true mappings and minimise false mappings throughout the text preprocessing pipeline. From a statistical perspective, it is a local optimisation for syntactic matching that will benefit the global optimisation for OM, as syntactic matching provides “anchor mappings” for lexical and semantic matching.
  3. (3)
    The use of LLMs is not a one-size-fits-all solution for OM tasks. OM is a complex multi-level matching mixed with syntactic, lexical, and semantic matching. LLMs may not outperform traditional methods in performing these sub-matching tasks. The development and implementation of an LLM-based framework that incorporates different matching techniques is a new research direction for OM, but it is beyond the scope of this paper. We refer readers to several recent works in [40, 41, 42, 43, 44].
@@
This study also demonstrates that the nature of text preprocessing used in NLP applications can be context-based. Not all pipeline methods are effective, and some of them may even hamper the overall performance of downstream tasks. Modifications are required before applying text preprocessing to specific NLP tasks.
VIII Limitations
@@
OAEI track repositories are comprehensive but do not cover all real-world scenarios. Furthermore, the reference files contained in the OAEI track repositories are not “complete” gold standards and some need further development. In this study, we only deal with equivalence mappings between named classes, object properties, and data type properties. Disjointness mappings are not considered. Finding widely used baseline data with ground truth matches for subsumptions is empirically difficult. Only a few OAEI track repositories contain subsumption mappings, and they are not enough to robustly demonstrate text preprocessing used in subsumption mappings. This work has not been entered into the official OAEI campaign because of its intense focus on only the preprocessing step.
IX Conclusion
@@
In this paper, we conduct a comprehensive study on the effect of the text preprocessing pipeline on syntactic OM. 8 OAEI track repositories with 49 distinct alignments are evaluated. Despite the importance of text preprocessing in syntactic OM, our experimental results indicate that the text preprocessing pipeline is currently ill-equipped to handle OM tasks. (1) We find that Phase 1 text preprocessing methods (Tokenisation and Normalisation) help with both matching completeness (i.e. Recall) and correctness (i.e. Precision). Phase 2 text preprocessing methods (Stop Words Removal and Stemming/Lemmatisation) are less effective. They can improve matching completeness (i.e. Recall), but matching correctness (i.e. Precision) is relatively low. (2) We propose a novel context-based pipeline repair approach to repair the less effective Phase 2 text preprocessing methods. By using a reserved word set to reduce false positive samples detected, our approach outperforms the traditional text preprocessing pipeline, in particular, the matching correctness (i.e. Precision) and overall matching performance (i.e. F1 Score). (3) We also argue the traditional text preprocessing pipeline is not obsolete in the era of LLMs, but an integral part for LLMs used in the OM tasks.
Our future work will focus on handling class axioms and complex relationships to evaluate the text preprocessing pipeline for OM tasks. We will also study the pipeline and pipeline repair approach working with both knowledge-based OM systems and LLM-based OM systems.
Supplemental Material Statement
The source code for this study is available from https://github.com/qzc438-research/ontology-nlp. The OAEI track repositories used in this study can be downloaded from OAEI MELT at https://dwslab.github.io/melt/track-repository (date accessed: 2024-06-01). The OAEI data policy can be found in https://oaei.ontologymatching.org/doc/oaei-deontology.2.html.
Acknowledgment
@@
The authors would like to thank Alice Richardson of the Statistical Support Network, Australian National University (ANU), for helpful advice on the statistical analysis in this paper. The authors also thank the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.

### v4

@@ 6.2. Experiment Setup
    We report our results over a single run. There could be slight differences in the results over multiple trails, but these differences are not significant with respect to the ranking and trend of different LLMs and prompt templates.
Figure 10. Discovery rate of TPs using LLMs on the Largebio Track.
Figure 11. The discovery rate of FPs using LLMs on the Largebio Track.
6.3. Detecting True Mappings
LLMs are expected to answer “yes” for TPs. The discovery rate is defined as the number of “yes” answers compared to the total number of TPs. For LLMs, we directly prompt LLMs to determine whether two entities are identical or not. For the text preprocessing pipeline, we employ the pipeline via function calling (a.k.a tool calling). This means that the LLMs act as an interface, and mapping judgement is based on programme logic defined in the text preprocessing method. Two entities are considered to be matched if they have the same results after the text preprocessing.
Figure 10 shows the discovery rate of TPs using LLMs on the Largebio Track. Here we use only Tokenisation (T) and Normalisation (N) to preprocess the text because Stop Words Removal (R) and Stemming/Lemmatisation (S/L) are shown to be less effective from our analysis in Section 4. Although some API-accessed commercial LLMs have shown the capability to detect TPs without a text preprocessing step, this behaviour is non-deterministic across different LLMs and prompt templates. For different LLMs, open-source LLMs perform better than commercial API-accessed LLMs. It could be problematic for some commercial API-accessed LLMs (e.g. claude-3haiku) to find two identical entities if text preprocessing is not applied prior to the comparison. They may consider entities with different formatting that are used in different contexts and therefore have different meanings. For different prompt templates, few-shot (PT2) significantly improves correctness, but self-reflection (PT3) may even hamper performance. The combination of few-shot and self-reflection (PT4) may even perform worse than the simple basic prompt (PT1). We believe that this phenomenon is mainly caused by LLM “overthinking”, a type of LLM hallucinations in which LLMs could over-analyse user input and generate unnecessarily complex details. We also need to note here that the example generated in the few-shot prompt is based on our observation of the existing TPs. Finding a good example in real-world applications is not easy because ontology entities can have different formatting by nature. On the other hand, using the classical text preprocessing pipeline via function calling could help LLMs produce a stable result with a high discovery rate (almost 100% from our observation) across different models and prompt templates.
6.4. Detecting False Mappings
LLMs are expected to answer “no” for FPs. The discovery rate is defined as the number of “no” answers compared to the total number of FPs. Specifically, these FPs are non-trivial mappings that the classical text preprocessing methods cannot detect.
Figure 11 shows the discovery rate of FPs using LLMs on the Largebio Track. Here we use LLMs to detect FPs generated by the Lancaster Stemmer (SL). SL has shown a lower precision compared to other text preprocessing methods, which is caused by the method producing a significant number of FPs. LLMs have shown a strong capability to discover the FPs produced by the classical text preprocessing pipeline. For different models, API-accessed commercial LLMs are better than open-source LLMs and reach a higher discovery rate of more than 80%. For different prompt templates, both few-shot (PT2), self-reflection (PT3), and a combination of the above (PT4) are better than a simple basic prompt (PT4). While mapping repair is usually non-trivial and requires domain expertise, LLMs have shown potential for performing a high-quality post hoc correction without human intervention.
6.5. Implications
LLMs were originally developed and trained for question answering (QA). OM is not only a QA task, but also a knowledge-intensive task that requires a comprehensive understanding and reasoning of domain knowledge. The studies in this section demonstrate the limitations and opportunities of integrating the text preprocessing pipeline with LLMs in syntactic OM. We highlight two pathways for this integration. One pathway is to use function calling to invoke the classical text preprocessing pipeline within LLMs, which packages the programming-based text preprocessing pipeline as a tool for LLMs to use. This provides stable results for text preprocessing in syntactic OM. The other pathway is to use LLMs as a repair tool after text preprocessing, which could be a cost-effective solution for post hoc correction. This can also be extended using human-in-the-loop to validate the repaired results generated by LLMs and feed the validated data back into LLMs so that they can better understand the context of the task (i.e. fine-tuning LLMs used in syntactic OM).
7. Discussion
OM aims to bridge the gap by capturing similar concepts that occur in different ontologies. OM is a complex process that combines syntactic matching, lexical matching, and semantic matching. Each level of matching may use different matching techniques, for example, text preprocessing for syntactic matching, text vectorisation for lexical matching, and logical reasoning for semantic matching. This study concentrates on a common practice using the text preprocessing pipeline in syntactic OM. While generic text preprocessing pipeline methods in syntactic OM are usually applied on the basis of intuition or extrapolation from other experiences, this work advances the state of knowledge towards making design decisions objective and supported by evidence.
@@ 7. Discussion
    Our study shows (a) whether to use, or not use, and (b) how to use these text preprocessing methods. Such experimental results will benefit decision-making when selecting appropriate text preprocessing methods for syntactic OM. For large-scale OM, it can significantly reduce unnecessary trial costs.
@@
    Our context-based pipeline repair approach is proposed to repair the less effective Phase 2 text preprocessing methods. Its broader value is showing how to maximise true mappings and minimise false mappings throughout the text preprocessing pipeline. From a statistical perspective, it is a local optimisation for syntactic matching that will benefit the global optimisation for OM, as syntactic matching provides “anchor mappings” for lexical matching and semantic matching.
@@
    The use of LLMs is not a one-size-fits-all solution for OM tasks. OM is a complex multi-level matching mixed with different matching techniques. LLMs may not outperform traditional methods in performing these sub-matching tasks. We suggest “rules of thumb” for syntactic OM. If two ontology entities can be matched using the intuitive text preprocessing pipeline, there is no need to leverage LLMs for the same purpose.
@@
    The development and implementation of an LLM-based framework that incorporates different matching techniques is a new research direction for OM, but it is beyond the scope of this paper. We refer the reader to several recent works in (He et al., 2023; Norouzi et al., 2023; Hertling and Paulheim, 2023; Qiang et al., 2023; Giglou et al., 2024; Amini et al., 2024).
@@
    This study also demonstrates that the nature of text preprocessing used in NLP applications can be context-based. Not all pipeline methods are effective, and some of them may even hamper the performance of downstream tasks. Modifications are required before applying text preprocessing to specific NLP tasks.
8. Limitations
@@
The references contained in the OAEI tracks are not “complete” gold standards and some need further development. In this study, we only deal with equivalence mappings between classes and properties. Subsumption mappings are not considered. Finding widely used baseline data with ground truth matches for subsumptions is empirically difficult. Only a few OAEI tracks contain subsumption mappings, and they are not enough to robustly demonstrate the text preprocessing used in subsumption mappings. This work has not been included in the official OAEI campaign because of its intense focus on only text preprocessing in syntactic OM.
9. Conclusion
@@
In this paper, we conduct a comprehensive study on the effect of the text preprocessing pipeline on syntactic OM. 8 OAEI tracks with 49 distinct alignments are evaluated. Despite the importance of text preprocessing in syntactic OM, our experimental results indicate that the text preprocessing pipeline is currently ill-equipped. (1) We find that Phase 1 text preprocessing methods (Tokenisation and Normalisation) help with both matching completeness (i.e. recall) and correctness (i.e. precision). Phase 2 text preprocessing methods (Stop Words Removal and Stemming/Lemmatisation) are less effective. They can improve matching completeness (i.e. recall), but matching correctness (i.e. precision) is relatively low. (2) We propose a novel context-based pipeline repair approach to repair the less effective Phase 2 text preprocessing methods. By using a reserved word set to reduce false positive samples detected, our approach outperforms the traditional text preprocessing pipeline, in particular, the matching correctness (i.e. precision) and overall matching performance (i.e. F1 score). (3) We also argue that the traditional text preprocessing pipeline is not obsolete in the era of LLMs. It is an integral part of the LLMs used in OM tasks. Integrating the classical text preprocessing pipeline can significantly improve the robustness of purely prompt-based approaches. Our future work will focus on handling class axioms and complex relationships to evaluate the text preprocessing pipeline. We will also study the pipeline and pipeline repair approach working with both knowledge-based OM systems and LLM-based OM systems.
Acknowledgements.
@@
The authors thank Sven Hertling for curating the Ontology Alignment Evaluation Initiative (OAEI) datasets. The authors thank Ernesto Jimenez-Ruiz for helpful advice on reproducing the OAEI Largebio Track. The authors thank Alice Richardson of the Statistical Support Network, Australian National University (ANU), for helpful advice on the statistical analysis in this paper. The authors thank the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.